# FlashI2V: Fourier-Guided Latent Shifting Prevents Conditional Image Leakage in Image-to-Video Generation

### v1

@@ Abstract
In Image-to-Video (I2V) generation, a video is created using an input image as the first-frame condition. Existing I2V methods concatenate the full information of the conditional image with noisy latents to achieve high fidelity. However, the denoisers in these methods tend to shortcut the conditional image, which is known as conditional image leakage, leading to performance degradation issues such as slow motion and color inconsistency. In this work, we further clarify that conditional image leakage leads to overfitting to in-domain data and decreases the performance in out-of-domain scenarios. Moreover, we introduce Fourier-Guided Latent Shifting I2V, named FlashI2V, to prevent conditional image leakage. Concretely, FlashI2V consists of: (1) Latent Shifting. We modify the source and target distributions of flow matching by subtracting the conditional image information from the noisy latents, thereby incorporating the condition implicitly. (2) Fourier Guidance. We use high-frequency magnitude features obtained by the Fourier Transform to accelerate convergence and enable the adjustment of detail levels in the generated video. Experimental results show that our method effectively overcomes conditional image leakage and achieves the best generalization and performance on out-of-domain data among various I2V paradigms. With only 1.3B parameters, FlashI2V achieves a dynamic degree score of 53.01 on Vbench-I2V, surpassing CogVideoX1.5-5B-I2V and Wan2.1-I2V-14B-480P. Github page: https://pku-yuangroup.github.io/FlashI2V/
@@ 1 Introduction
Conditional Video Generation (Ren et al., 2024; Zhang et al., 2023; Hu, 2024; Li et al., 2025a; Yuan et al., 2025b; Yu et al., 2024; Zhang et al., 2025) refers to the technology that generates videos based on user-provided conditions, with significant applications being Text-to-Video (T2V) Generation and Image-to-Video (I2V) Generation. Since T2V generation produces a video solely based on a prompt, it struggles to accurately define scenes, such as accurate color and shape within the video. In contrast, I2V generation creates a video from both a user-provided image and a descriptive prompt, ensuring that the video content semantically aligns with the prompt and the first frame matches the provided image at the pixel level. In the commercial State-of-the-Art (SOTA) video generation product, Kling (AI, 2025), 85% of usage calls are for I2V generation.
Leveraged in I2V methods including Stable Video Diffusion (SVD) (Blattmann et al., 2023), Open-Sora Plan (Lin et al., 2024a), CogVideoX (Yang et al., 2024), and Wan2.1 (Wan et al., 2025), existing approaches concatenate the conditional image latents encoded by a Variational Autoencoder (VAE) (Kingma & Welling, 2013) with the noisy latents along the channel dimension, achieving exceptionally high fidelity for the first frame. However, previous works (Zhao et al., 2024; Choi et al., 2025) highlight that existing methods suffer from conditional image leakage. Especially at large time steps, the denoiser directly utilizes the condition in a shortcut manner to minimize loss instead of performing the complex denoising process during training, resulting in slow motion in the generated output during inference. In addition to slow motion, we also observe other performance degradation issues such as color inconsistency in the generated video, as shown in Fig. 1(a).
@@
To investigate why conditional image leakage leads to performance degradation, we explore the generalization of the existing concatenating I2V paradigm. During training, the conditional image is the first frame of a video. In contrast, during inference, the conditional image can come from any source and is not necessarily the first frame of an existing video. The ability to generate reasonable and high-quality videos from any conditional image requires strong generalization in I2V methods. Since we cannot achieve the training dataset of any existing model, we train a model with weights initialized from Wan-T2V-1.3B using the existing concatenating I2V paradigm and compare its performance on both in-domain and out-of-domain data. Each video is divided into temporal chunks with an equal frame interval. We then compare the Fréchet Video Distance (FVD) (Unterthiner et al., 2018) of the generated chunks with the ground truth chunks to assess the generation quality at different time points in the video. Theoretically, the first frame of the generated video must exactly match the conditional image, while subsequent frames lack such constraints, resulting in an increasing chunk-wise FVD over time. In a desired I2V paradigm, this increasing pattern should hold for both in-domain and out-of-domain data. As illustrated in Fig. 1(b), experimental results reveal that chunk-wise FVD on in-domain data increases gradually over time. However, in out-of-domain data, chunk-wise FVD remains consistently high. By comparing the chunk-wise FVD variation patterns on in-domain and out-of-domain data, we conclude that even if the first frame matches the conditional image, shortcutting causes out-of-domain results to lack coherent video quality. The law learned from in-domain data fails to generalize to out-of-domain data, indicating that the concatenating paradigm faces an overfitting challenge, and a more reasonable paradigm is expected.
@@
Compared to various existing I2V paradigms, only FlashI2V demonstrates the same FVD variation pattern on both in-domain and out-of-domain data, indicating that it avoids the leakage caused by shortcutting the conditional image. Furthermore, FlashI2V achieves the lowest FVD value among different I2V paradigms, showcasing its excellent performance. With only 1.3B parameters, FlashI2V achieves comparable scores to CogVideoX1.5-5B-I2V and Wan2.1-I2V-14B-480P on Vbench-I2V (Huang et al., 2024; Zheng et al., 2025) and obtains a dynamic degree score of 53.01, significantly outperforming the other two methods with larger parameter sizes.
@@ 2.1 Text-to-Video Generation
In recent years, Text-to-Video Generation has made significant progress. T2V methods often use diffusion models (Ho et al., 2020; Song et al., 2020a; b) to model the generation process. Previous works typically employ UNet (Ronneberger et al., 2015) and add temporal transformers after image weights (commonly referred to as the 2+1D paradigm) as denoisers (Yuan et al., 2025c; 2024; Guo et al., 2023; Wang et al., 2025; Chen et al., 2023a; 2024a). After the release of Sora (Brooks et al., 2024), the community uses Diffusion Transformers (DiTs) (Peebles & Xie, 2023; Yao et al., 2025) as denoisers (Zheng et al., 2024; Lin et al., 2024a; Ma et al., 2024; Xu et al., 2024) within the 2+1D paradigm to achieve T2V. To overcome the limited capability of the 2+1D paradigm in temporal modeling, approaches like Open-Sora Plan v1.2 (Lin et al., 2024a) model all tokens uniformly (commonly referred to as the 3D paradigm) instead of differentiating between image and temporal weights. At present, with the adoption of 3D Transformer and more advanced diffusion models, flow matching (Lipman et al., 2022; Liu, 2022), T2V models can generate highly realistic videos.
@@ 2.2 Image-to-Video Generation
Image-to-Video Generation leverages a conditional image and a prompt as inputs, enhancing the controllability of the generated video. Stable Video Diffusion (SVD) (Blattmann et al., 2023) combines conditional image latents with noisy latents, and injects high-level semantic information of the conditional image extracted by CLIP (Radford et al., 2021; Zhu et al., 2023; Lin et al., 2023; 2024b; Chen et al., 2024b) into the denoiser. DynamiCrafter (Xing et al., 2024) improves on SVD by using a query transformer (Li et al., 2023) to extract CLIP tokens. CogVideoX (Yang et al., 2024) concatenates zero-padding conditional image latents with noisy latents to introduce conditions. SEINE (Chen et al., 2023b) introduces a temporal inpainting model, where conditional image latents and mask sequences are concatenated with noisy latents to fill in subsequent frames. Open-Sora Plan v1.3 (Lin et al., 2024a; Li et al., 2025b) further expands the inpainting model to cover more tasks and proposes a progressive training strategy to enhance performance. Wan2.1 (Wan et al., 2025) improves the inpainting model by introducing semantic information extracted by CLIP. All of these approaches inject full conditional image information into the denoiser through a concatenation operation, resulting in excellent fidelity for the first frame.
@@ 2.3 Conditional Image Leakage
Conditional image leakage (Zhao et al., 2024; Yuan et al., 2025a) is an issue where the model shortcuts the conditional image information, especially at large time steps, rather than utilizing it as an auxiliary to generate the video from noisy latents. SVD introduces adding a small amount of noise to the conditional image to increase the dynamic degree, marking the first attempt to reduce conditional image leakage. Previous work (Zhao et al., 2024) proposes addressing the leakage by starting the generation process from an earlier time step during inference and designing a time-dependent noise distribution for the conditional image during training. Additionally, Adaptive Low-pass Guidance (ALG) (Choi et al., 2025) is a training-free approach by using the low-pass information of the conditional image rather than its full information at large time steps. At present, resolving conditional image leakage remains an open problem, with no universally accepted solution in the community.
@@ 3.1 Preliminary for Flow Matching
Continuous Normalizing Flows (CNFs) (Chen, 2018) aim to learn a transformation from a sample $\bm{z}_{1}$ from a source distribution $q_{1}(\bm{z})$ to a sample $\bm{z}_{0}$ from a target distribution $q_{0}(\bm{z})$, where $q_{0}(\bm{z})$ represents the data distribution, and $q_{1}(\bm{z})$ is typically a known prior distribution, such as a standard normal distribution. This transformation is usually modeled as an ordinary differential equation (ODE) with $t\in[0,1]$. Let $\bm{z}_{t}$ represents the intermediate state from $\bm{z}_{1}$ to $\bm{z}_{0}$, then the transformation is governed by the following equation:
@@
Here, $\bm{v}_{t}(\bm{z}_{t},t)$ defines the velocity field at any time, dictating how the distribution transfers over time. The concept of Flow Matching (FM) (Lipman et al., 2022; Liu, 2022; Tong et al., 2023) is to directly learn the vector field $\bm{v}_{t}(\bm{z}_{t},t)$ from $\bm{z}_{1}$ to $\bm{z}_{0}$ using a neural network $\bm{v}_{\theta}(\bm{z}_{t},t)$. Specifically, for $\bm{z}_{1}\sim q_{1}$ and $\bm{z}_{0}\sim q_{0}$, their linear interpolation is constructed as follows:
@@
where $\mathcal{U}[0,1]$ represents the uniform distribution over $[0,1]$. During sampling, we sample $\bm{z}_{1}\sim q_{1}$ and solve the ODE $\frac{d\bm{z}_{t}}{dt}=\bm{v}_{\theta}(\bm{z}_{t},t,\bm{y})$ step by step from $t=1$ to $t=0$ to obtain $\bm{z}_{0}\sim q_{0}$. Unlike Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), which require the source distribution to be a standard normal distribution, FM imposes no such constraints on the source distribution. This flexibility allows FM to be used for transferring between any two distributions.
@@ 3.2 Latent Shifting
We consider implementing I2V without explicitly incorporating the full information of the conditional image into the hidden states of the denoiser. Let $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$, where $\mathcal{N}(\mathbf{0},{\bm{I}})$ denotes a standard normal distribution. Let a conditional image be $\bm{S}\in\mathbb{R}^{c\times h\times w}$, and a video starting with the conditional image be $\bm{X}\in\mathbb{R}^{c\times t\times h\times w}$, which means $\bm{X}[:,0]=\bm{S}$. Let $\mathcal{E}$ represent the encoder of the Variational Autoencoder (VAE) (Kingma & Welling, 2013). Denote the source distribution sample as $\bm{z}_{1}^{T}$ and the target distribution sample as $\bm{z}_{0}^{T}$ for the T2V task, we have $\bm{z}_{1}^{T}=\bm{\epsilon}$, $\bm{z}_{0}^{T}=\bm{x}=\mathcal{E}(\bm{X})$, and the intermediate state $\bm{z}_{t}^{T}$ at any time $t$ under FM is given by:
@@
$\bm{z}_{t}^{I}=(1-t)\bm{z}_{1}^{I}+t\bm{z}_{0}^{I}=\beta\bm{z}_{t}^{T}+[\alpha+(\gamma-\alpha)t]\bm{s}+(\kappa-\beta)t\bm{x}.$ (8)
@@
$\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\frac{d\bm{z}_{t}^{I}}{dt}=\beta\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)+(\gamma-\alpha)\bm{s}+(\kappa-\beta)\bm{x}.$ (9)
When training an I2V model, we typically inherit the weight of the corresponding T2V model. A good initialization can leverage knowledge from the pre-trained weights as much as possible. It is observed that when $\gamma=\alpha$ and $\kappa=\beta=1$, we have $\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)$, meaning the optimization objectives for I2V and T2V are structurally the same. In this case, $\bm{z}_{t}^{I}$ can be expressed as:
$\bm{z}_{t}^{I}=\bm{z}_{t}^{T}+\alpha\bm{s}.$ (10)
When $\alpha=0$, we have $\bm{z}_{t}^{I}=\bm{z}_{t}^{T}$, meaning that without the conditional image as a condition, the model is equivalent to a T2V model. When $\alpha\neq 0$, the input of the denoiser incorporates conditional image information. In this case, we have $\bm{z}_{1}^{I}=\bm{\epsilon}-(-\alpha\bm{s})$ and $\bm{z}_{0}^{I}=\bm{x}-(-\alpha\bm{s})$, where $\alpha\bm{s}$ can be viewed as the shifting of the latents in I2V task relative to T2V task and we aim to learn the conditional image information from the shifted latents.
Furthermore, $\alpha$ can be viewed as a constant weight for $\bm{s}$. Since the effective information of each position varies, we can replace $\alpha\bm{s}$ with $\bm{\phi}(\bm{s})$, where $\bm{\phi}(\cdot)$ is a learnable projection. Additionally, we can zero-initialize $\bm{\phi}(\cdot)$, ensuring that the input distribution of the denoiser is not disrupted at the start of training. Since $\bm{\phi}(\cdot)$ is a network that is independent of time $t$, we still have $\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)$. Now we obtain the ultimate form for the I2V method based on the latent shifting:

### v2

@@ Abstract
In Image-to-Video (I2V) generation, a video is created using an input image as the first-frame condition. Existing I2V methods concatenate the full information of the conditional image with noisy latents to achieve high fidelity. However, the denoisers in these methods tend to shortcut the conditional image, which is known as conditional image leakage, leading to performance degradation issues such as slow motion and color inconsistency. In this work, we further clarify that conditional image leakage leads to overfitting to in-domain data and decreases the performance in out-of-domain scenarios. Moreover, we introduce Fourier-Guided Latent Shifting I2V, named FlashI2V, to prevent conditional image leakage. Concretely, FlashI2V consists of: (1) Latent Shifting. We modify the source and target distributions of flow matching by subtracting the conditional image information from the noisy latents, thereby incorporating the condition implicitly. (2) Fourier Guidance. We use high-frequency magnitude features obtained by the Fourier Transform to accelerate convergence and enable the adjustment of detail levels in the generated video. Experimental results show that our method effectively overcomes conditional image leakage and achieves the best generalization and performance on out-of-domain data among various I2V paradigms. With only 1.3B parameters, FlashI2V achieves a dynamic degree score of 53.01 on Vbench-I2V, surpassing CogVideoX1.5-5B-I2V and Wan2.1-I2V-14B-480P. Project page: https://pku-yuangroup.github.io/FlashI2V/
@@ 1 Introduction
Conditional Video Generation (consisti2v; i2vgen_xl; animate_anyone; anyi2v; consisid; viewcrafter; fantasyid) refers to the technology that generates videos based on user-provided conditions, with significant applications being Text-to-Video (T2V) Generation and Image-to-Video (I2V) Generation. Since T2V generation produces a video solely based on a prompt, it struggles to accurately define scenes, such as accurate color and shape within the video. In contrast, I2V generation creates a video from both a user-provided image and a descriptive prompt, ensuring that the video content semantically aligns with the prompt and the first frame matches the provided image at the pixel level. In the commercial State-of-the-Art (SOTA) video generation product, Kling (kling), 85% of usage calls are for I2V generation.
Leveraged in I2V methods including Stable Video Diffusion (SVD) (stable_video_diffusion), Open-Sora Plan (open_sora_plan), CogVideoX (cogvideox), and Wan2.1 (wan), existing approaches concatenate the conditional image latents encoded by a Variational Autoencoder (VAE) (vae) with the noisy latents along the channel dimension, achieving exceptionally high fidelity for the first frame. However, previous works (conditional_image_leakage; ALG) highlight that existing methods suffer from conditional image leakage. Especially at large time steps, the denoiser directly utilizes the condition in a shortcut manner to minimize loss instead of performing the complex denoising process during training, resulting in slow motion in the generated output during inference. In addition to slow motion, we also observe other performance degradation issues such as color inconsistency in the generated video, as shown in Fig. 1(a).
@@
To investigate why conditional image leakage leads to performance degradation, we explore the generalization of the existing concatenating I2V paradigm. During training, the conditional image is the first frame of a video. In contrast, during inference, the conditional image can come from any source and is not necessarily the first frame of an existing video. The ability to generate reasonable and high-quality videos from any conditional image requires strong generalization in I2V methods. Since we cannot achieve the training dataset of any existing model, we train a model with weights initialized from Wan-T2V-1.3B using the existing concatenating I2V paradigm and compare its performance on both in-domain and out-of-domain data. Each video is divided into temporal chunks with an equal frame interval. We then compare the Fréchet Video Distance (FVD) (fvd) of the generated chunks with the ground truth chunks to assess the generation quality at different time points in the video. Theoretically, the first frame of the generated video must exactly match the conditional image, while subsequent frames lack such constraints, resulting in an increasing chunk-wise FVD over time. In a desired I2V paradigm, this increasing pattern should hold for both in-domain and out-of-domain data. As illustrated in Fig. 1(b), experimental results reveal that chunk-wise FVD on in-domain data increases gradually over time. However, in out-of-domain data, chunk-wise FVD remains consistently high. By comparing the chunk-wise FVD variation patterns on in-domain and out-of-domain data, we conclude that even if the first frame matches the conditional image, shortcutting causes out-of-domain results to lack coherent video quality. The law learned from in-domain data fails to generalize to out-of-domain data, indicating that the concatenating paradigm faces an overfitting challenge, and a more reasonable paradigm is expected.
@@
Compared to various existing I2V paradigms, only FlashI2V demonstrates the same FVD variation pattern on both in-domain and out-of-domain data, indicating that it avoids the leakage caused by shortcutting the conditional image. Furthermore, FlashI2V achieves the lowest FVD value among different I2V paradigms, showcasing its excellent performance. With only 1.3B parameters, FlashI2V achieves comparable scores to CogVideoX1.5-5B-I2V and Wan2.1-I2V-14B-480P on Vbench-I2V (vbench; vbench2) and obtains a dynamic degree score of 53.01, significantly outperforming the other two methods with larger parameter sizes.
@@ 2.1 Text-to-Video Generation
In recent years, Text-to-Video Generation has made significant progress. T2V methods often use diffusion models (ddpm; ddim; score_diffusion) to model the generation process. Previous works typically employ UNet (unet) and add temporal transformers after image weights (commonly referred to as the 2+1D paradigm) as denoisers (magictime; chronomagic; animatediff; lavie; videocrafter1; videocrafter2). After the release of Sora (sora), the community uses Diffusion Transformers (DiTs) (dit; lightingdit) as denoisers (open_sora; open_sora_plan; latte; easyanimate) within the 2+1D paradigm to achieve T2V. To overcome the limited capability of the 2+1D paradigm in temporal modeling, approaches like Open-Sora Plan v1.2 (open_sora_plan) model all tokens uniformly (commonly referred to as the 3D paradigm) instead of differentiating between image and temporal weights. At present, with the adoption of 3D Transformer and more advanced diffusion models, flow matching (flow_matching; rectified_flow), T2V models can generate highly realistic videos.
@@ 2.2 Image-to-Video Generation
Image-to-Video Generation leverages a conditional image and a prompt as inputs, enhancing the controllability of the generated video. Stable Video Diffusion (SVD) (stable_video_diffusion) combines conditional image latents with noisy latents, and injects high-level semantic information of the conditional image extracted by CLIP (clip; languagebind; lin2023video; lin2024moe; chen2024sharegpt4video) into the denoiser. DynamiCrafter (dynamicrafter) improves on SVD by using a query transformer (query_transformer) to extract CLIP tokens. CogVideoX (cogvideox) concatenates zero-padding conditional image latents with noisy latents to introduce conditions. SEINE (seine) introduces a temporal inpainting model, where conditional image latents and mask sequences are concatenated with noisy latents to fill in subsequent frames. Open-Sora Plan v1.3 (open_sora_plan; li2025wf) further expands the inpainting model to cover more tasks and proposes a progressive training strategy to enhance performance. Wan2.1 (wan) improves the inpainting model by introducing semantic information extracted by CLIP. All of these approaches inject full conditional image information into the denoiser through a concatenation operation, resulting in excellent fidelity for the first frame.
@@ 2.3 Conditional Image Leakage
Conditional image leakage (conditional_image_leakage; opens2v) is an issue where the model shortcuts the conditional image information, especially at large time steps, rather than utilizing it as an auxiliary to generate the video from noisy latents. SVD introduces adding a small amount of noise to the conditional image to increase the dynamic degree, marking the first attempt to reduce conditional image leakage. Previous work (conditional_image_leakage) proposes addressing the leakage by starting the generation process from an earlier time step during inference and designing a time-dependent noise distribution for the conditional image during training. Additionally, Adaptive Low-pass Guidance (ALG) (ALG) is a training-free approach by using the low-pass information of the conditional image rather than its full information at large time steps. At present, resolving conditional image leakage remains an open problem, with no universally accepted solution in the community.
@@ 3.1 Preliminary for Flow Matching
Continuous Normalizing Flows (CNFs) (CNFs) aim to learn a transformation from a sample $\bm{z}_{1}$ from a source distribution $q_{1}(\bm{z})$ to a sample $\bm{z}_{0}$ from a target distribution $q_{0}(\bm{z})$, where $q_{0}(\bm{z})$ represents the data distribution, and $q_{1}(\bm{z})$ is typically a known prior distribution, such as a standard normal distribution. This transformation is usually modeled as an ordinary differential equation (ODE) with $t\in[0,1]$. Let $\bm{z}_{t}$ represents the intermediate state from $\bm{z}_{1}$ to $\bm{z}_{0}$, then the transformation is governed by the following equation:
@@
Here, $\bm{v}_{t}(\bm{z}_{t},t)$ defines the velocity field at any time, dictating how the distribution transfers over time. The concept of Flow Matching (FM) (flow_matching; rectified_flow; improving_flow_matching) is to directly learn the vector field $\bm{v}_{t}(\bm{z}_{t},t)$ from $\bm{z}_{1}$ to $\bm{z}_{0}$ using a neural network $\bm{v}_{\theta}(\bm{z}_{t},t)$. Specifically, for $\bm{z}_{1}\sim q_{1}$ and $\bm{z}_{0}\sim q_{0}$, their linear interpolation is constructed as follows:
@@
where $\mathcal{U}[0,1]$ represents the uniform distribution over $[0,1]$. During sampling, we sample $\bm{z}_{1}\sim q_{1}$ and solve the ODE $\frac{d\bm{z}_{t}}{dt}=\bm{v}_{\theta}(\bm{z}_{t},t,\bm{y})$ step by step from $t=1$ to $t=0$ to obtain $\bm{z}_{0}\sim q_{0}$. Unlike Denoising Diffusion Probabilistic Models (DDPM) (ddpm), which require the source distribution to be a standard normal distribution, FM imposes no such constraints on the source distribution. This flexibility allows FM to be used for transferring between any two distributions.
@@ 3.2 Latent Shifting
We consider implementing I2V without explicitly incorporating the full information of the conditional image into the hidden states of the denoiser. Let $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$, where $\mathcal{N}(\mathbf{0},{\bm{I}})$ denotes a standard normal distribution. Let a conditional image be $\bm{S}\in\mathbb{R}^{c\times h\times w}$, and a video starting with the conditional image be $\bm{X}\in\mathbb{R}^{c\times t\times h\times w}$, which means $\bm{X}[:,0]=\bm{S}$. Let $\mathcal{E}$ represent the encoder of the Variational Autoencoder (VAE) (vae). Denote the source distribution sample as $\bm{z}_{1}^{T}$ and the target distribution sample as $\bm{z}_{0}^{T}$ for the T2V task, we have $\bm{z}_{1}^{T}=\bm{\epsilon}$, $\bm{z}_{0}^{T}=\bm{x}=\mathcal{E}(\bm{X})$, and the intermediate state $\bm{z}_{t}^{T}$ at any time $t$ under FM is given by:
@@
$\bm{z}_{t}^{I}=(1-t)\bm{z}_{0}^{I}+t\bm{z}_{1}^{I}=\kappa\bm{z}_{t}^{T}+[\gamma+(\alpha-\gamma)t]\bm{s}+(\beta-\kappa)t\bm{\epsilon}.$ (8)
@@
$\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\frac{d\bm{z}_{t}^{I}}{dt}=\kappa\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)+(\alpha-\gamma)\bm{s}+(\beta-\kappa)\bm{\epsilon}.$ (9)
When training an I2V model, we typically inherit the weight of the corresponding T2V model. A good initialization can leverage knowledge from the pre-trained weights as much as possible. It is observed that when $\alpha=\gamma$ and $\beta=\kappa=1$, we have $\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)$, meaning the optimization objectives for I2V and T2V are structurally the same. In this case, $\bm{z}_{t}^{I}$ can be expressed as:
$\bm{z}_{t}^{I}=\bm{z}_{t}^{T}+\gamma\bm{s}.$ (10)
When $\gamma=0$, we have $\bm{z}_{t}^{I}=\bm{z}_{t}^{T}$, meaning that without the conditional image as a condition, the model is equivalent to a T2V model. When $\gamma\neq 0$, the input of the denoiser incorporates conditional image information. In this case, we have $\bm{z}_{1}^{I}=\bm{\epsilon}-(-\gamma\bm{s})$ and $\bm{z}_{0}^{I}=\bm{x}-(-\gamma\bm{s})$, where $\gamma\bm{s}$ can be viewed as the shifting of the latents in I2V task relative to T2V task and we aim to learn the conditional image information from the shifted latents.
Furthermore, $\gamma$ can be viewed as a constant weight for $\bm{s}$. Since the effective information of each position varies, we can replace $\gamma\bm{s}$ with $\bm{\phi}(\bm{s})$, where $\bm{\phi}(\cdot)$ is a learnable projection. Additionally, we can zero-initialize $\bm{\phi}(\cdot)$, ensuring that the input distribution of the denoiser is not disrupted at the start of training. Since $\bm{\phi}(\cdot)$ is a network that is independent of time $t$, we still have $\bm{v}_{t}^{I}(\bm{z}_{t}^{I},t)=\bm{v}_{t}^{T}(\bm{z}_{t}^{T},t)$. Now we obtain the ultimate form for the I2V method based on the latent shifting: